OpenStudy (anonymous):

Suppose you have a drawer full of white, black, and yellow pairs of socks. If the probability of picking a white pair of socks is 4/9, and the probability of picking a black pair of socks is 7/18 , what is the probability of picking a yellow pair of socks?

OpenStudy (anonymous):

a. 1/6 b 5/25 c 7/15 d 16/27

OpenStudy (alekos):

well i reckon there's 18 socks in total, so that would mean 8 white, 7 black and therefore 3 yellow. so the P(y) = 3/18 = 1/6

OpenStudy (anonymous):

thnk you alot really